Skip to Search

 
diagonal spacer

Systems Staff Meeting Notes
March 5, 2007

Present: Thomas Bielicke, Cynthia Bowker, Will Chaney, Logan Clapp, Debra Fetzer, Winston Harris, Bobbie Meng, Tom Minton, Chris Nicolich, Amy Polk, Mark Sullivan, Stephen Williams
Announcements
  1. Bobbie's mom celebrated her 87th birthday. Carlton gave her crown molding as a gift.
  2. Stephen had to get vitamins and emergency food for his chinchilla - he is dangerously underweight. The remedy is working and now he is eating.
  3. Duct cleaning to begin this week and weekend.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Staff and DLC workstations - on hold until we can determine the size of the monitors. We will set up a demo to test 17", 19", 21" and dual set-up monitors with ALEPH v. 18 (set to the correct resolution) and come to a consensus as to what size monitor should be purchased.
  2. Quote from Dell for the replacement of NT2R - 6 terrabytes. Would have a duplicate unit at MSL.
  3. We are also looking at virtual servers to consolidate.
  4. Michelle Foss is having trouble with Dreamweaver - it has not yet been reported to SysHelp.
Liaisons
  1. Liaison training started today. It's going well - the classes are full.
  2. ALEPH problems to be reported through SysHelp? Yes.
Networking
  1. Daylight Savings Time patches - the Exchanged server was updated and we pushed them to the mailboxes this weekend. There still may be some errors. The client patch was pushed earlier. We still need to patch the W2K boxes. Anything you find wrong now needs to be fixed manually. These patches are supposed to fix the beginning and ending date for DST.
  2. West security videos - nothing new.
  3. Xerox - nothing. Brett has been out-of-town.
  4. VOIP at MSL - reports of drops. CNS has done an IIOS upgrade and are scheduled to upgrade the switches in two weeks.
  5. NT9 relocation - it has been moved down to Systems, but the table is still in the old location. It needs to be returned to Systems. We have not reconfigured it yet due to the dead tape drive. After shipping confusion, the new drive arrived and has been installed. Discussion of MAC address.
  6. Campus AD - down the road.
Software Projects Status
  1. ERES Security and Carrel Reservation Systems are both ongoing.
  2. Winston changed the Business Services invoices to meet ILL requirements.
  3. Last Wednesday, the version of ILLIAD changed. Logan helped with it. Minor problems with Web pages.
  4. NT18 - failed to do the Patron Updates again. It was out of space on the E: drive. Winston cleaned it up. Should be fine now.
  5. Mark has set up training for Tom in DLC to provide better understanding of the project.
  6. Training database - log-in button is nice. HR wants monthly statistics generated.
Web Projects
  1. Campus-wide Web Management Committee Meeting set for Thursday, March 29 at 1PM. Feel free to come if you are interested. Reitz Union Auditorium.
Systems Security Management
  1. Meeting held.

 

University of Florida Home Page